How do I use log_a(20)=1.4406 and log_a(5)=0.7740 to evaluate each expression.
1. log_a(100)
2. log_a 5^3

Answer by stanbon(75887)

How do I use log_a(20)=1.4406 and log_a(5)=0.7740 to evaluate each expression.
1. log_a(100)
100 = 5*20
--------------------
Ans: loga(100) = loga(5) + loga(20) = 0.7740+1.4406 = 2.2146
----------------------
2. log_a 5^3 = 3*loga(5) = 3*0.7740 = 3.3220
------------------------
Cheers,
Stan H.
----------------